OpenStudy (dinorap1):

How do I solve this system using the elimination method? PLEASE HELP! 5x+4y=12 and 3x-3y=18

OpenStudy (triciaal):

hint the second equation can be factored

OpenStudy (dinorap1):

What's happening here? 3x−3y=18 >x−y=6> 4x−4y=24

OpenStudy (dinorap1):

Oh, did you delete your response...?

OpenStudy (triciaal):

the objective of the elimination method is to pick a variable and make the coefficients equal and opposite so that will be eliminated when the equivalent equations are added

OpenStudy (solomonzelman):

\(5x+4y=12\) \(3x-3y=18\) > \(x-y=6\) > \(4x-4y=24\) divided times by 6 6 \(5x+4y=12\) \(4x-4y=24\) ~~~~~~~~~ Add the two equations now, and tell me what you get.

OpenStudy (solomonzelman):

I deleted my reply, to modify it....

OpenStudy (solomonzelman):

MISTAKE: Oh, divide by 3, not 6.

OpenStudy (triciaal):

correction equation 2 divided by 3

OpenStudy (solomonzelman):

yes, just said that

OpenStudy (triciaal):

mails crossed

OpenStudy (solomonzelman):

:)

OpenStudy (dinorap1):

Oh, I see what you're doing now. Don't you mean multiplied by 4 for the 2nd step?

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

Then you have to add the equations.

OpenStudy (dinorap1):

I was taught a different method to solve this though. Are you sure this is a valid way to solve for the elimination method?

OpenStudy (solomonzelman):

yes, I am positive

OpenStudy (solomonzelman):

(good that you noticed my typo errors, thank you)

OpenStudy (triciaal):

glad you are paying attention

OpenStudy (triciaal):

he choose to get rid of y's now add and solve for x

OpenStudy (triciaal):

yes that is the way and I was never in any of his classes

OpenStudy (dinorap1):

Alright I'll solve for it then! 9x=36 > (9/9)x=(36/9) > x=4

OpenStudy (solomonzelman):

yes, x=4 is right.

OpenStudy (dinorap1):

And then you plug the value of x into one of the equations to get y. I can understand it from here! Thank you two so much for the help! :)

OpenStudy (solomonzelman):

sure, yw

OpenStudy (triciaal):

no problem